OpenStudy (usarmy3947):

Which expression is equivalent to 7-2?

OpenStudy (highschoolmom2010):

are there choices?

OpenStudy (usarmy3947):

7 × -2 -7 × 2 × 2 1/ 7.2.2 1/7.7

OpenStudy (anonymous):

Not clear ... Which one is power or ...??? x or * ???

OpenStudy (usarmy3947):



OpenStudy (phi):

if it is \( 7^{-2} \) you can write that as 7^(-2) with an uparrow ^ to show (-2) is an exponent the -2 means that you "flip" the 7: \[ \frac{1}{7^2} \] 7^2 means 7*7 so the answer is \[ \frac{1}{49} \]

OpenStudy (anonymous):

can you write you answers better please ??:)

OpenStudy (usarmy3947):

ah thats not a answer choice

OpenStudy (usarmy3947):

Simplify .

OpenStudy (anonymous):

I would have to say 7x - 2

OpenStudy (anonymous):

2 * 2 * 2 = 2 * 2 = 4. 4 * 2 = 8

OpenStudy (usarmy3947):

cool